Handwrite in text plz not in pic since its hard to read from and plz dont copy answers that were answered before least 2-3 paragraphs

1. Your friend has just purchased a house and has incurred a $150,000, 4.5% mortgage payable at $760.03 per month. After making the first monthly payment, he receives a statement from the bank indicating only $197.53 had been applied to reducing the principal amount of the loan. Your friend then calculates that at the rate of $197.53 per month, it will take 63 years to pay off the $150,000 mortgage. Discuss and explain whether your friend’s analysis is correct or not.

'2.Discuss and explain why a company may choose to raise capital by issuing bonds instead of issuing stock

least 2-3 paragraphs
Answer #1

1.

No, Friend analysis is wrong. Because every month contribution to capital will increase and interest payable will decrease.

While paying 1st month EMI interest portion will be $562.50((150,000*4.5%)*1/12), and principal is $197.53

While Paying 2nd month EMI Interest portion will be $561.76((150,000-197.53)*4.5%)*1/12 and principal is $198.27

So, every month from EMI contribution towards principal increases.

Based on information given.

EMI = P × r × (1 + r)n/((1 + r)n - 1) where P= Loan amount, r= interest rate, n=tenure in number of months

760.03=150,000*0.045*(1+.045)n/(((1+0.045)n-1))

n = 360 months which is equal to 30 years. Loan will finish in 30 years

2. Reasons to raise capital by issuing bonds instead of issuing stock.

By issuing bonds you will fix your cost of capital to certain percentage that will be mostly less than your return on capital. So, the future gains from capital raised from Bonds will be contributing to Earning per share which in turn increase Stock value, company value etc.

issuing new capital will increase no of investors and ownership will increase proportionally. So in turn it will decrease earning per share. Decrease in EPS will reduce the firm value.

Amount paid as interest on bond will be considered for tax deductions. Which in turn will have again positive impact on EPS.

When the existing group of shareholders does not want to have their ownership interests come down by the sale of shares to new investors, they will push for a bond issuance. Since bonds are a form of debt, no new shares will be sold.(only in case of Non - Convertable bonds)